OpenStudy (anonymous):

A trip from Ohio to New York is 529 miles. What equation shows the time t it takes to go by car if r is the average speed during the trip?

OpenStudy (anonymous):

So let the distance covered be d. Now total time taken = t. And we know that avg speed(r) = Total distance(d)/Total time(t) r = d/t t = d/r t = 529/r , where d = 529 miles distance

OpenStudy (anonymous):

thanks!